OpenStudy (maddiie):

(5,1) is a solution to the equation 2x-3y=7. What does this mean?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

the point is given as (x,y). So x=5 and y=1. If that point is a solution to the equation when we substitue x=5 and y=1 in the equation will be true. We have: 2(5)-3(1)=10-3=7, 7=7, the equation is true so (5,1) is a solution.

OpenStudy (maddiie):

Thanks this is my first time using this so i don't knowwhere the button Good answer is at.

OpenStudy (anonymous):

it's in line with my name on the upper right hand side of my post i think